Appliance Center is an experienced home appliance dealer. Appliance Center also offers a number of services for the home appliances that it sells. Assume that Appliance Center sells ovens on a standalone basis. Appliance Center also sells installation services and maintenance services for ovens. However, Appliance Center does not offer installation or maintenance services to customers who buy ovens from other vendors. Pricing for ovens is as follows.

Oven only $800
Oven with installation service 850
Oven with maintenance services 1,000
Oven with installation and maintenance services 990


In each instance in which maintenance services are provided, the maintenance service is separately priced within the arrangement at $200. Additionally, the incremental amount charged by Appliance Center for installation approximates the amount charged by independent third parties. Ovens are sold subject to a general right of return. If a customer purchases an oven with installation and/or maintenance services, in the event Appliance Center does not complete the service satisfactorily, the customer is only entitled to a refund of the portion of the fee that exceeds $800.

Assume that a customer purchases an oven with both installation and maintenance services for $990. Based on its experience, Appliance Center believes that it is probable that the installation of the equipment will be performed satisfactorily to the customer. Assume that the maintenance services are priced separately (i.e., the three components are distinct).

(b) Indicate the amount of revenue that should be allocated to the oven, the installation, and to the maintenance contract.
Answer #1

Individual cost for each item:

Oven = 800

Installation = 850-800=50

Maintenance= 200

Total : 1,050

Allocation: cost of individual item * amount paid /total cost of individual item

Oven : 800*990/1,050=$ 754.29

Installation : 50*990/1,050=$ 47.14

Maintenance : 200*990/1,050=$188.57

I HOPE IT USEFUL TO YOU IF YOU HAVE ANY DOUBT PLZ COMMENT GIVE ME UP-THUMB. THANKS....